# Design of Bayesian Clinical Trials with Clustered Data and Multiple Endpoints

### v2

@@ Abstract
In the design of clinical trials, it is essential to assess the design operating characteristics (i.e., the probabilities of making correct decisions). Common practice for the evaluation of operating characteristics in Bayesian clinical trials relies on estimating the sampling distribution of posterior summaries via Monte Carlo simulation. It is computationally intensive to repeat this estimation process for each design configuration considered, particularly for clustered data that are analyzed using complex, high-dimensional models. In this paper, we propose an efficient method to assess operating characteristics and determine sample sizes for Bayesian trials with clustered data and multiple endpoints. We prove theoretical results that enable posterior probabilities to be modeled as a function of the sample size. Using these functions, we assess operating characteristics at a range of sample sizes given simulations conducted at only two sample sizes. These theoretical results are also leveraged to quantify the impact of simulation variability on our sample size recommendations. The applicability of our methodology is illustrated using a current cluster-randomized Bayesian adaptive clinical trial with multiple endpoints.
@@ 1 Introduction
Bayesian methods for data-driven decision making have become increasingly popular in the design and analysis of clinical trials. Bayesian inference is particularly useful in trials with clustered data given the challenges with numerical stability that can arise in frequentist generalized linear mixed modeling (Breslow, 2004; Capanu et al., 2013). Decision making in Bayesian clinical trials is often based on posterior and posterior predictive probabilities (Spiegelhalter et al., 2004; Berry et al., 2010). In many cases, despite the use of Bayesian analysis and decision criteria, grant and regulatory agencies require the design to be assessed with respect to frequentist operating characteristics (FDA, 2019). Certain operating characteristics, such as power and the type I error rate, are considered for most Bayesian trials. Additional operating characteristics, including the probabilities of making certain interim decisions, are also of interest in adaptive designs. In Bayesian trials, sampling distributions of posterior summaries must be accurately estimated to correctly assess the frequentist operating characteristics. Sample size determination (SSD) procedures rely on accurate estimates of these operating characteristics.
@@
Wang and Gelfand (2002) proposed a general framework for Bayesian SSD in which Monte Carlo simulation is used to accurately estimate frequentist operating characteristics of Bayesian designs. This computational approach estimates sampling distributions by simulating many repetitions of a trial with a particular set of design, analysis, and decision parameters. In comprehensive trial design, these numerical studies can involve many scenarios, each of which is defined by a different combination of design, analysis, and decision parameters. Design parameters may include sample sizes in adaptive and non-adaptive contexts; analysis parameters may include the effect size for the estimand of interest and values for nuisance parameters; and decision parameters may include thresholds for efficacy or futility. These simulations often borrow from the frequentist framework for power analysis in that each scenario and its corresponding sampling distribution are defined given fixed choices for the analysis model parameters. Alternative Bayesian SSD approaches in which sampling distributions of posterior summaries are marginalized with respect to a design prior that incorporates uncertainty in the model parameters have also been proposed (O’Hagan et al., 2005; De Santis, 2007; Gubbiotti and De Santis, 2011).
@@
The U.S. Food and Drug Administration (FDA) recommends using at least $10^{4}$ simulation repetitions to estimate the sampling distribution for each scenario considered (FDA, 2019). Each simulation repetition for each scenario typically involves computational posterior approximation. The computational burden associated with posterior approximation is substantial for high-dimensional models, including those for clustered data that use random effects to model within-cluster dependence. Clustered data are common in clinical settings (e.g., longitudinal studies and cluster-randomized trials). In trials where adjusted generalized linear models are employed for analysis, further computing resources are required to construct marginal estimands via Bayesian G-computation (Kalton, 1968; Keil et al., 2018; Daniel et al., 2021; Willard et al., 2024). These estimands are obtained by marginalizing samples from the posterior distribution with respect to the distribution of additional covariates or random effects. Drawing inference based on marginal estimands is important when population average treatment effects are of interest but adjusted analyses are employed (Willard et al., 2024).
Moreover, many modern designs leverage co-primary or secondary outcomes in decision making, resulting in complex decision criteria. Each of these endpoints addresses a distinct aspect of a disease, and it may be necessary to show the treatment provides clinical benefit across all or certain combinations of the outcomes (FDA, 2022). The literature on multiple endpoints in clinical trials primarily focuses on correcting for multiplicity in traditional designs (see e.g., Holm (1979); Dmitrienko et al. (2008); Bretz et al. (2009)), with minimal consideration of Bayesian approaches. Since accounting for multiple endpoints further complicates computational posterior approximation, it is important to develop methodology that ensures numerical studies efficiently inform operating characteristic assessment and Bayesian SSD. Streamlined design methods reduce costs associated with trial design and enhance timely communication between the stakeholders of a clinical trial.
@@
Various strategies have recently been proposed to reduce the computational burden associated with estimating operating characteristics for Bayesian designs. Golchi (2022) proposed a modeling approach to estimate sampling distributions of posterior summaries and trial design operating characteristics using Gaussian processes; Golchi and Willard (2024) presented an alternative modeling method that augmented simulations with asymptotic theory. Hagar and Stevens (2024) approximated power curves using segments of the relevant sampling distributions. For a given simulation scenario, Hagar and Stevens (2025) proposed a method to assess operating characteristics across the sample size space using estimates of the sampling distribution at only two sample sizes. Hagar et al. (2025) extended that method to accommodate sequential designs. However, all the aforementioned work focuses on a single endpoint informed by independent observations. Since decisions in many trials are based on multiple estimands for dependent data, we build upon the method from Hagar and Stevens (2025) to facilitate the efficient assessment of operating characteristics in designs with clustered data and complex decision rules. We also illustrate how to accommodate adaptive designs with clustered data by combining the theory from this paper with results from Hagar et al. (2025). Our proposed methods are simple to implement and promote an economical framework for simulation-based Bayesian SSD that is suitable for broad use with clinical trials.
@@
The remainder of this article is structured as follows. In Section 2, we introduce preliminary concepts required to describe our methods. In Section 3, we prove new theoretical results about a proxy to the sampling distribution of posterior probabilities based on clustered data. In Section 4, we adapt these theoretical results to develop a procedure for Bayesian SSD that requires estimation of the sampling distribution of posterior probabilities at only two sample sizes. This procedure also allows users to efficiently quantify the impact of simulation variability on the sample size recommendation. Section 5 applies our method to a Bayesian adaptive trial where households are randomized to preventative tuberculosis treatments and interim decisions are based on two secondary outcomes and a primary safety endpoint. This trial serves as a substantive application that motivates our SSD approach, which can be broadly implemented to design Bayesian trials. We conclude with a summary and discussion of extensions to this work in Section 6.
2 Preliminaries
2.1 Operating Characteristics with Multiple Estimands
Consider a clinical trial where multiple outcomes related to treatment efficacy and safety must be evaluated. Each of these $K\geq 1$ outcomes is summarized by a scalar estimand, giving rise to a vector of $K$ estimands. Decisions in Bayesian trials are made using predefined decision rules based on posterior summaries of the estimands (Berry et al., 2010). This paper focuses on trials where decision criteria are defined with respect to posterior probabilities. Because the main contribution of this paper addresses clustered designs with multiple endpoints, we introduce our notation and methods in a non-adaptive setting. Please see Hagar et al. (2025) for a discussion of adaptive designs. The statistical model for the trial is defined using a set of parameters $\boldsymbol{\theta}\in\boldsymbol{\Theta}$. The estimands are a function of these parameters: $\boldsymbol{\delta}(\boldsymbol{\theta})\in\mathbb{R}^{K}$. The hypotheses that inform decision making are formulated as
$H_{0}:\boldsymbol{\delta}(\boldsymbol{\theta})\notin(\boldsymbol{\delta}_{L},\boldsymbol{\delta}_{U})~{}~{}~{}\text{vs.}~{}~{}~{}H_{1}:\boldsymbol{\delta}(\boldsymbol{\theta})\in(\boldsymbol{\delta}_{L},\boldsymbol{\delta}_{U}),$ (1)
where $\boldsymbol{\delta}_{L}$ and $\boldsymbol{\delta}_{U}$ represent vectors of $K$ interval endpoints. This general notation for $\boldsymbol{\delta}_{L}$ and $\boldsymbol{\delta}_{U}$ accommodates hypothesis tests based on superiority, noninferiority, and practical equivalence (Spiegelhalter et al., 1994, 2004). The $k^{\text{th}}$ row of (1) outlines complementary hypotheses for the scalar estimand $\delta_{k}(\boldsymbol{\theta})$, defined as $H_{0,k}:\delta_{k}(\boldsymbol{\theta})\notin(\delta_{L,k},\delta_{U,k})$ vs. $H_{1,k}:\delta_{k}(\boldsymbol{\theta})\in(\delta_{L,k},\delta_{U,k})$. Using this notation, $H_{1}$ is true if and only if all subhypotheses $\{H_{1,K}\}_{k=1}^{K}$ are true. Otherwise, $H_{0}$ is true.
@@ 2.1 Operating Characteristics with Multiple Estimands
Multi-component decision criteria are defined using combinations of posterior probabilities about the individual estimands $\{\delta_{k}(\boldsymbol{\theta})\}_{k=1}^{K}$ in $\boldsymbol{\delta}(\boldsymbol{\theta})$. The posterior of $\boldsymbol{\theta}$ synthesizes information from the prior distribution for $\boldsymbol{\theta}$ and the data $\mathcal{D}_{c}$. Since this paper focuses on clustered data, $\mathcal{D}_{c}$ consists of $N$ observations from clusters $j=1,\dots,c$. There are $n_{j}$ observations from cluster $j$ such that $\sum_{j=1}^{c}n_{j}=N$. As described later, we index the data $\mathcal{D}_{c}$ and consider SSD in terms of the number of clusters $c$. This framework accommodates various contexts where clustered data arise. In cluster-randomized trials, individual observations correspond to distinct participants that are randomized to a given treatment in groups or clusters (Murray, 1998). Cluster-randomized trials are practical in many real world settings and reduce the risk of contamination between treatment groups (Hemming et al., 2021). In longitudinal studies, individual observations correspond to repeated measurements from the same participant.
The data $\mathcal{D}_{c}=\{{\bf{Y}}_{N\times K},{\bf{A}}_{N},{\bf{X}}_{N\times p}\}$ may consist of observed outcomes for each estimand ${\bf{Y}}_{N\times K}$, treatment assignments ${\bf{A}}_{N}$, and $p$ additional covariates ${\bf{X}}_{N\times p}$. While we use this general notation, certain outcomes may not be related to all $p$ additional covariates. These data are used to compute a vector of $K$ posterior probabilities regarding the subhypotheses $\{H_{1,k}\}_{k=1}^{K}$:
$\boldsymbol{\tau}(\mathcal{D}_{c})=\begin{bmatrix}\tau_{1}(\mathcal{D}_{c})\\ \tau_{2}(\mathcal{D}_{c})\\ \vdots\\ \tau_{K}(\mathcal{D}_{c})\end{bmatrix}=\begin{bmatrix}\Pr(H_{1,1}~{}|~{}\mathcal{D}_{c})\\ \Pr(H_{1,2}~{}|~{}\mathcal{D}_{c})\\ \vdots\\ \Pr(H_{1,K}~{}|~{}\mathcal{D}_{c})\end{bmatrix}.$ (2)
Decision rules are defined by comparing $\boldsymbol{\tau}(\mathcal{D}_{c})$ to decision thresholds $\boldsymbol{\gamma}\in[0,1]^{K}$. Depending on the context, the decision threshold $\boldsymbol{\gamma}$ could guide conclusions about treatment superiority or futility. Complex decision rules in clinical trials may be constructed by considering any subset of the $K$ individual comparisons between $\boldsymbol{\tau}(\mathcal{D}_{c})$ and $\boldsymbol{\gamma}$. To accommodate a broad range of decision rules, we define the binary indicator $\nu(\mathcal{D}_{c})$ that equals 1 if and only if the decision criteria are met when comparing $\boldsymbol{\tau}(\mathcal{D}_{c})$ with $\boldsymbol{\gamma}$.

### v3

@@ Abstract
In the design of clinical trials, it is essential to assess the design operating characteristics (e.g., power and the type I error rate). Common practice for the evaluation of operating characteristics in Bayesian clinical trials relies on estimating the sampling distribution of posterior summaries via Monte Carlo simulation. It is computationally intensive to repeat this estimation process for each design configuration considered, particularly for clustered data that are analyzed using complex, high-dimensional models. In this paper, we propose an efficient method to assess operating characteristics and determine sample sizes for Bayesian trials with clustered data. We prove theoretical results that enable posterior probabilities to be modeled as a function of the number of clusters. Using these functions, we assess operating characteristics at a range of sample sizes given simulations conducted at only two cluster counts. These theoretical results are also leveraged to quantify the impact of simulation variability on our sample size recommendations. The applicability of our methodology is illustrated using an example cluster-randomized Bayesian clinical trial.
@@ 1 Introduction
Bayesian methods for data-driven decision making have become increasingly popular in the design and analysis of clinical trials. Bayesian inference is particularly useful in trials with clustered data given the challenges with numerical stability that can arise in frequentist generalized linear mixed modeling (Breslow, 2004; Capanu et al., 2013). Decision making in Bayesian clinical trials is often based on posterior and posterior predictive probabilities (Spiegelhalter et al., 2004; Berry et al., 2010). In many cases, despite the use of Bayesian analysis and decision criteria, grant and regulatory agencies require the design to be assessed with respect to frequentist operating characteristics (FDA, 2019). Certain operating characteristics, such as power and the type I error rate, are considered for most Bayesian trials. In Bayesian trials, sampling distributions of posterior summaries must be accurately estimated to correctly assess the frequentist operating characteristics. Sample size determination (SSD) procedures rely on accurate estimates of these operating characteristics.
@@
Wang and Gelfand (Wang and Gelfand, 2002) proposed a general framework for Bayesian SSD in which Monte Carlo simulation is used to accurately estimate frequentist operating characteristics of Bayesian designs. This computational approach estimates sampling distributions by simulating many repetitions of a trial with a particular set of design, analysis, and decision parameters. In comprehensive trial design, these numerical studies can involve many scenarios, each of which is defined by a different combination of design, analysis, and decision parameters. Design parameters may include sample sizes; analysis parameters may include the effect size for the estimand of interest and values for nuisance parameters; and decision parameters may include decision thresholds. These simulations often borrow from the frequentist framework for power analysis in that each scenario and its corresponding sampling distribution are defined given fixed choices for the analysis model parameters. Alternative Bayesian SSD approaches in which sampling distributions of posterior summaries are marginalized with respect to a design prior that incorporates uncertainty in the model parameters have also been proposed (O’Hagan et al., 2005; De Santis, 2007; Gubbiotti and De Santis, 2011).
@@
The U.S. Food and Drug Administration (FDA) recommends using at least $10^{4}$ simulation repetitions to estimate the sampling distribution for each scenario considered (FDA, 2019). Each simulation repetition for each scenario typically involves computational posterior approximation. The computational burden associated with posterior approximation is substantial for high-dimensional models, including those for clustered data that use random effects to model within-cluster dependence. Clustered data are common in clinical settings (e.g., longitudinal studies and cluster-randomized trials).
In cluster-randomized trials, cluster-level or individual-level estimands may be of interest depending on the research objectives (Kahan et al., 2023). This paper is concerned with individual-level estimands. In trials where adjusted generalized linear models are employed for analysis, further computing resources are required to construct marginal estimands via Bayesian G-computation (Kalton, 1968; Keil et al., 2018; Daniel et al., 2021; Willard et al., 2024). These estimands are obtained by marginalizing samples from the posterior distribution with respect to the distribution of additional covariates or random effects. Drawing inference based on marginal estimands is important when population average treatment effects are of interest but adjusted analyses are employed (Willard et al., 2024).
@@
It is important to develop methodology that ensures numerical studies efficiently inform operating characteristic assessment and Bayesian SSD. Streamlined design methods reduce costs associated with trial design and enhance timely communication between the stakeholders of a clinical trial. Various strategies have recently been proposed to reduce the computational burden associated with estimating operating characteristics for Bayesian designs. Golchi (2022) proposed a modeling approach to estimate sampling distributions of posterior summaries and trial design operating characteristics using Gaussian processes; Golchi and Willard (2024) presented an alternative modeling method that augmented simulations with asymptotic theory. Hagar and Stevens (2024) approximated power curves using segments of the relevant sampling distributions.
For a given simulation scenario, Hagar and Stevens (2025) proposed a method to assess operating characteristics across the sample size space using estimates of the sampling distribution at only two sample sizes. In that work, the quantiles of the sampling distribution of the logits of posterior probabilities were shown to depend linearly on the sample size. Hence, numerical studies at only two sample sizes are required to infer the sampling distribution at new sample sizes. However, that contribution and the other aforementioned work focus on Bayesian studies with independent observations. Since decisions in many trials are based on dependent data, we build upon the method from Hagar and Stevens (2025) to facilitate the efficient assessment of operating characteristics in designs with clustered data. Our proposed methods are simple to implement and promote an economical framework for simulation-based Bayesian SSD that is suitable for broad use with clinical trials.
@@
The remainder of this article is structured as follows. In Section 2, we describe a Bayesian trial where households are randomized to preventative tuberculosis treatments. This illustrative example contextualizes our paper’s methodological development. We introduce background information and notation required to describe our methods in Section 3. In Section 4, we prove new theoretical results about a proxy to the sampling distribution of posterior probabilities based on clustered data. In Section 5, we adapt these theoretical results to develop a procedure for Bayesian SSD that requires estimation of the sampling distribution of posterior probabilities at only two cluster counts. In Section 6, we conduct numerical studies to assess the performance of the proposed SSD method for our illustrative example. Section 7 discusses extensions to this work for adaptive trials and trials with multiple endpoints. We conclude with a discussion in Section 8.
2 Example, Data, and Estimands
In this section, we introduce general notation for clustered data in clinical trials. This structure for the data is contextualized using an illustrative example. This example draws inspiration from the Shorter and Safer Treatment Regimens for Latent Tuberculosis (SSTARLET) trial (Menzies, 2024). SSTARLET aims to establish that new tuberculosis preventive medications are non-inferior to a reference treatment with respect to safety. It is a cluster-randomized adaptive platform trial in which all members of a household are assigned to the same treatment arm, and new regimens are compared against a common control arm across multiple endpoints. To emphasize the applicability of our methodology with a straightforward yet representative trial design, our illustrative example simplifies SSTARLET to a parallel, non-adaptive design in which a single new treatment is compared with the reference arm using one endpoint.
@@ 2 Example, Data, and Estimands
We now introduce general notation for the design of trials with clustered data. We denote the data to be collected in the trial by $\mathcal{D}_{c}$, consisting of $N$ observations from clusters $j=1,\dots,c$. There are $n_{j}$ observations from cluster $j$ such that $\sum_{j=1}^{c}n_{j}=N$. As described later, we index the data $\mathcal{D}_{c}$ and consider SSD in terms of the number of clusters $c$. This framework accommodates various contexts where clustered data arise. In cluster-randomized trials, individual observations correspond to distinct participants that are randomized to a given treatment in groups or clusters (Murray, 1998). In longitudinal studies, individual observations correspond to repeated measurements from the same participant.
The data $\mathcal{D}_{c}=\{{\bf{Y}}_{N},{\bf{A}}_{N},{\bf{X}}_{N\times p}\}$ may consist of observed outcomes ${\bf{Y}}_{N}$, treatment assignments ${\bf{A}}_{N}$, and $p$ additional covariates ${\bf{X}}_{N\times p}$. For clustered data, $q$ unobserved random effects ${\bf{W}}_{c\times q}$ account for the dependence between outcomes from the same cluster. We assume that these random effects are not shared between observations from different clusters. As a result, we make the standard assumption that outcomes in different clusters are independent.
To illustrate this notation within our example, the outcomes ${\bf{Y}}_{N}$ indicate whether or not participants experienced a grade 3 to 5 adverse event (AE). There are binary treatment indicators ${\bf{A}}_{N}$ but no additional covariates ${\bf{X}}_{N\times p}$ for this example. We assume the outcomes in this trial come from a Bayesian logistic regression model with cluster-specific intercepts ${\bf{W}}_{c\times 1}$ as random effects. This model is
$y_{j,i}\sim\text{BIN}(1,\text{expit}(\eta_{j,i})),~~\text{where}~~\eta_{j,i}=\beta_{0}+\beta_{1}A_{j,i}+w_{j},$
where the subscripts are such that $j$ indexes the cluster and $i$ indexes the observation within the $j^{\text{th}}$ cluster. The $\beta_{0}$ and $\beta_{1}$ parameters represent a common intercept and slope and $w_{j}\sim\mathcal{N}(0,\sigma^{2})$ are cluster-specific intercepts. Using general notation, the statistical model for the trial is defined using a set of parameters $\boldsymbol{\theta}\in\boldsymbol{\Theta}$ that may include the distributional parameters for the random effects. The parameters for our example are $\boldsymbol{\theta}=(\beta_{0},\beta_{1},\sigma^{2})$.
The data associated with the outcomes ${\bf{Y}}_{N}$ are summarized by a scalar estimand. As mentioned earlier, we are interested in the marginal individual-level treatment effect. To facilitate the introduction of our methodology, we denote the estimand as a function of the model parameters: $\delta(\boldsymbol{\theta})\in\mathbb{R}$. This paper focuses on marginal estimands $\delta(\boldsymbol{\theta})$ that can be expressed as a contrast of population quantities for each treatment. We let $A\in\{0,1\}$ denote the treatment assignment, where $A=1$ and $A=0$ respectively represent being randomized to the treatment and control groups. We let $\mu(\boldsymbol{\theta};A)$ be the population mean, marginalized with respect to the distributions of $\bf{X}$} and $\bf{W}$}, corresponding to the treatment group $A$:
$\mu(\boldsymbol{\theta};A)=\int_{{\bf{X}}}\int_{{\bf{W}}}\mu(\boldsymbol{\theta};A,{\bf{X}},{\bf{W}})p({\bf{X}})p({\bf{W}})d{\bf{W}}d{\bf{X}}.$ (1)
Using a difference-based contrast without loss of generality, we broadly represent the marginal estimand of interest as
$\delta(\boldsymbol{\theta})=\mu(\boldsymbol{\theta};A=1)-\mu(\boldsymbol{\theta};A=0).$ (2)
For illustration, the estimand for our example is the difference between the marginal AE rates in the treatment and control arms:
$\delta(\boldsymbol{\theta})=\int_{w}\text{expit}(\beta_{0}+\beta_{1}+w)p(w)dw-\int_{w}\text{expit}(\beta_{0}+w)p(w)dw.$ (3)
Lastly, we discuss why our methodology provides the option to marginalize over the distribution of the random effects $\bf{W}$}. Most cluster-specific models are predicated on the existence of latent risk groups that are indexed by the cluster-specific effects (Neuhaus et al., 1991). These latent risk groups may not be balanced for all treatment groups in trials with clustered data, impacting the validity of the trial’s conclusions. Since these cluster-specific effects are not observed, it is difficult to determine whether they are balanced between treatment groups. We therefore focus on the case where inference is conducted by treating $\bf{W}$} as additional variables that we marginalize over. Nevertheless, the theory and methods we develop here can also be applied to trials with conditional estimands $\delta(\boldsymbol{\theta};{\bf{X}},{\bf{W}})$ since conditional estimands are conceptually simpler than marginal ones in regression settings.
3 Computational Bottlenecks in Trial Design
We now discuss the computational burden associated with assessing design operating characteristics in clinical trials. Design operating characteristics are defined with respect to hypotheses about the estimand(s) of interest. For trials with a single primary estimand, the hypotheses that inform decision making are formulated as
$H_{0}:\delta(\boldsymbol{\theta})\notin(\delta_{L},\delta_{U})~~~\text{vs.}~~~H_{1}:\delta(\boldsymbol{\theta})\in(\delta_{L},\delta_{U}),$ (4)
where $\delta_{L}$ and $\delta_{U}$ represent lower and upper interval endpoints. This general notation for $\delta_{L}$ and $\delta_{U}$ accommodates hypothesis tests based on superiority, non-inferiority, and practical equivalence (Spiegelhalter et al., 1994, 2004). In this paper, we focus on trials where the decision rule is defined using the posterior probability that $H_{1}$ is true.
The posterior of $\boldsymbol{\theta}$ synthesizes information from the prior distribution for $\boldsymbol{\theta}$ and the data $\mathcal{D}_{c}$. In settings with random effects, Bayesian analysis is facilitated using the joint posterior of $(\boldsymbol{\theta},{\bf{W}})$ given the data, $\mathcal{D}_{c}$. Samples from the posterior distribution of $(\boldsymbol{\theta},{\bf{W}})$ give rise to posterior samples from the conditional parameter $\mu(\boldsymbol{\theta};A,{\bf{X}},{\bf{W}})$ with respect to fixed patterns for the covariates $\bf{X}$} and random effects $\bf{W}$}. Because the estimands in clinical trials are often non-collapsible (Daniel et al., 2021; Willard et al., 2024), we marginalize the posterior samples of $\mu(\boldsymbol{\theta};A,{\bf{X}},{\bf{W}})$ with respect to the distributions of $\bf{X}$} and $\bf{W}$} to obtain samples from the posterior distribution of $\mu(\boldsymbol{\theta};A)$ in (1). The procedure to implement this marginalization is based on reweighting according to repeated sampling from Dirichlet distributions; its details are described in Appendix B of the online supplement. We use this procedure to approximate the posterior distribution of $\delta(\boldsymbol{\theta})$. From this distribution, we can compute the following posterior probability:
$\tau(\mathcal{D}_{c})=\Pr(H_{1}~|~\mathcal{D}_{c}).$ (5)
In this work, we focus on decision rules that reject the null hypothesis if $\tau(\mathcal{D}_{c})\geq\gamma$, where $\gamma\in[0,1]$ is a decision threshold. The hypotheses in (4) are formulated such that we support the goal of the trial (e.g., superiority, non-inferiority, or practical equivalence) by rejecting $H_{0}$. For our illustrative example, the treatment arm is non-inferior to the control if the difference in marginal AE rates defined in (3) is less than 4%. To support the trial’s goal of non-inferiority, we therefore define the alternative hypothesis as $H_{1}:\delta(\boldsymbol{\theta})<0.04$. That is, $H_{1}$ is based on an interval such that $\delta_{L}=$-$\infty$ and $\delta_{U}=0.04$.